# The Radius Distribution of M dwarf-hosted Planets and its Evolution

### v1

@@ 4.1 Features of the M dwarf planet radius distribution
Distributions of planet radius with orbital period and stellar irradiance show the absence of (sub)Neptunes on short-period or high-irradiance orbits around solar-type stars, the so-called “Neptune desert" (Szabó & Kiss, 2011; Mazeh et al., 2016). There are no universally accepted boundaries for this desert but it is expected to depend on stellar mass (McDonald et al., 2019). Adopting the conservative criteria of a period of 2 days or 300, there is a single object in the Neptune desert around M dwarfs: KOI-952.05 (Kepler 32-f), a member of a multi-transiting planet system (Swift et al., 2013) with an orbital period of 0.74 days, has an estimated radius of , consistent with but more precise than that determined by Morton et al. (2016) and Berger et al. (2018) but significantly larger than previous estimates (e.g., Swift et al., 2013; Mann et al., 2013). Kepler-32f would be a relevant target to measure mass and search for atmospheric escape, except that it is faint ().
Caveats and limitations of our analysis
@@
We restricted our sample to planet host stars with detected rotational signals in Kepler lightcurves and these are a biased subset of the overall population. Younger stars will be rapidly rotating, magnetically active, and highly spotted and their rotational periods can be readily measured. However, due to rotational variability and intense flaring, it can be difficult to detect transiting planets (Gaidos et al., 2017). On the other hand, the oldest stars will be slowly rotating, less spotted, and exhibit less rotational variability, and the rotation periods can be similar to the Kepler quarterly roll ( days) and thus more difficult to retrieve (Santos et al., 2019). This is particularly true for the coolest stars in the sample (mid M-dwarfs) because these stars have the longest rotation periods for a given age. These selection effects can manifest themselves in the overall distribution of M dwarf planet host ages (e.g., Gaidos et al., 2023b). However, we see no trend of planet properties (i.e., radius) with host star properties (i.e., luminosity, Fig. 7).
@@
Transit parameters and stellar parameters were not computed self-consistently (the former assuming DR25 values and the latter based on revised luminosities and new empirical relations). Stellar density is covariant with transit impact parameter which, along with (which controls limb darkening coefficients) is covariant with the planet-to-star radius ratio . Lastly, our filtering of binaries from the Kepler late K- and early M-dwarf target sample was less thorough since AO imaging is not available for the overall Kepler K/M dwarf population as it is for the KOIs.
4.2 Clues to the nature of sub-Neptunes
The decline of the sub-Neptune population relative to super-Earths between the younger (3.8 Gyr) and older halves of our sample of late K and early M dwarf-hosted planets (Fig. 2), is analogous to what has been found among solar-type Kepler host stars (Berger et al., 2020; Sandoval et al., 2021; David et al., 2021; Chen et al., 2022). Berger et al. (2020) found a decline in the sub-Neptune peak in stars older than fiducial age of the Hyades (650 Myr) relative to younger stars, and interpreted this evolution in terms of the escape of H/He-rich envelopes. The Gyr timescale implied was seen as more supportive of the core-powered mechanism for mass-loss (Gupta & Schlichting, 2021) over photo-evaporative escape (Owen, 2019). The decline in sub-Neptunes between the younger and older halves of our sample of late-type host stars suggest a multi-Gyr timescale, and perhaps the same mechanism in operation around these lower-mass stars.
Figure 8 compares the period-radius distribution of the younger (red points) and older (blue points) halves of the sample. While there are both younger and older sub-Neptunes (, above the horizontal line) at periods days, nearly all those at longer periods (24 days) are in the younger half of the sample. The phenomenon manifests itself in a plot of stellar irradiance vs. radius (Fig. 9). The fraction of older sub-Neptunes in the sample falls from 54% at irradiances to 15% at lower irradiance (see red and blue histograms). There are no sub-Neptunes in the sample below an irradiance of 1 , although this could be due in part to low detection efficiency at long orbital periods. The two planets that are marginally less irradiated are KOIs 812.03 (Kepler 235-e) and 952.03 (Kepler 32-d), with ages of and Gyr, respectively (Gaidos et al., 2023b).
The disappearance of sub-Neptunes over several Gyr could be explained by thermal contraction of atmospheres combined with escape of H and He to space (Howe & Burrows, 2015, e.g.,). This evolution ultimately leaves super-Earth-sized cores, which on wider ( day) orbits are unlikely to be detected (grey contours in Fig. 8). However, this would not explain the relative abundance of older sub-Neptunes on closer orbits, where irradiation is higher and the process of evaporation and escape should be more rapid. One possibility is that the apparent dependence of sub-Neptune evolution on orbital period and irradiance is a selection effect, i.e. a combination of correlations of age with host star mass/luminosity, and host star mass and planet period distribution. There is, however no evidence for either correlation (Figs. 6 and 11).
The persistence of “warm" sub-Neptunes at older ages could be due to an inflation mechanism that is depends on the planet’s irradiation or equilibrium temperature and hence proximity to the host star. Turbet et al. (2020) showed that (primarily) rocky planets with water-dominated runaway greenhouse atmospheres could have sub-Neptune-like radii. Both Kopparapu et al. (2013) and Turbet et al. (2019) found the irradiance threshold for a runaway greenhouse on an Earth-like planet to be about 1.05 the present terrestrial value , while atmospheres containing significant H/He will have a lower threshold (Innes et al., 2023). Around M dwarfs stars the threshold is slightly lower because of reduced Rayleigh scattering and increased absorption by atmospheric H2O at the longer wavelengths where cooler stars emit more radiation; the value is predicted to be 0.9 for =3900K (M0 spectral type) (red line in Fig. 9). Three-dimensional general circulation models put the threshold somewhat higher and dependent on rotation rate (Leconte et al., 2013; Kodama et al., 2018, ; vertical orange dashed line in Fig. 9).
A significantly elevated runaway threshold could be produced by a much higher elevated Bond albedo of a cloudy steam atmosphere; that of Venus is around 0.76 (Moroz et al., 1985) but measurements of the phase function suggest a value as high as 0.9 (Mallama et al., 2006). This would move the runaway threshold to , the approximate boundary where sub-Neptunes become persistent (vertical dashed magenta line in Fig. 9). 666The albedo depends would depend strongly on cloud droplet size and the high reflectivity of Venus is largely due to sulfuric acid aerosols and a similar scenario might not apply to steam atmospheres.
Figure 9 includes predicted radius vs. irradiance for planets with 3 rocky cores and H2O mass fractions between 0.1 and 25% (Turbet et al., 2020), with the core mass chosen to reproduce the observed 1.5 population. Water mass fractions 25% would be needed to explain some objects. Water-rich planets have been invoked to explain the sub-Neptune population with radii up to 3 (Burn et al., 2024), but Vivien et al. (2022) found that pure water/steam atmospheres on sub-Earths become unstable if the atmospheric scale height exceeds 0.1 times the planet radius. The radii of some sub-Neptunes might also require significant H/He to explain, especially those of planets at larger separations; H/He lowers the mean molecular weight and decreases the irradiance threshold for a runaway water-vapor greenhouse (Innes et al., 2023). The rate of H/He escape would then control the (Gyr) timescale over which sub-Neptunes on orbits beyond the pure-water vapor runaway threshold shrink and became undetectable by Kepler. Escape of H/He from hotter sub-Neptunes inside the threshold might be slowed by dilution and cooling by H2O vapor (Yoshida et al., 2022), or counterbalanced by permanent magma oceans that act as a sink for oxygen released by UV photolysis of H2O and thus a source for H (Schaefer et al., 2016).
We investigated whether irradiance-dependent evolution of sub-Neptunes also appears in the sample of Kepler transiting planets around solar-type host stars with isochrone ages determined by Berger et al. (2018). We restrict the sample to those with ages more precise than 30%, and although the median age of the sample is 5 Gyr, we use the same division of 3.86 Gyr between young and old populations for comparison. No such effect is obvious in the distribution of planets with irradiance (Fig. 12). The ratio of young to old planets is statistically indistinguishable () for planets experiencing and . Three explanations for the difference between the late K/early M dwarf sample and solar-type stars are (1) the former is affected by small-number statistics and is a spurious result; (2) the ages of the Berger et al. (2018) sub-sample with relatively cool planets are more uncertain because the host stars are likely to be late G and early K dwarfs which evolve very slowly; and (3) there is a fundamental difference in the two populations of planets around these stars. Planets at a given bolometric irradiance around M dwarfs will experience far more X-ray and UV irradiation than their solar-type counterparts (McDonald et al., 2019) and it is this radiation that drives photoevaporation of H/He atmospheres (Owen, 2019). Thus temperate sub-Neptunes around solar-type stars may be dominated by cores with H/He atmospheres, while those around M dwarfs have lost most or all H/He and their radii are controlled by the behavior of water-dominated atmospheres. Measurements of planet mass by RVs and observations of atmospheres by JWST for examples around much closer, brighter host stars could distinguish between these scenarios.
There are six planets in the sample with irradiances less than and could be considered to orbit within the “habitable zones" of their host stars, but only two of these are super-Earth-sized: KOIs 571.05 (Kepler-186f Quintana et al., 2014), with a radius of and a gyrochronologic system age of Gyr, and KOI-2418.01 (Kepler-1229b, Morton et al., 2016), with a radius of and a gyrochronologic system age of Gyr.

### v2

@@ 3.1 Planet Radius-Period Distribution
As was shown for solar-type host stars (Fulton et al., 2017) and subsequently for late K- and M-type dwarfs (Hirano et al., 2018; Cloutier & Menou, 2020; Van Eylen et al., 2021), the planet distribution of this sample contains two distinct clusters at about 1.3 (“super-Earths") and 2.2 (“sub-Neptunes"), with an intervening gap or “valley" at 1.6-1.7. Around solar-type hosts there is a deficit of (sub-)Neptunes on short–period (or high–irradiance) orbits, the so-called “Neptune desert" (Szabó & Kiss, 2011; Mazeh et al., 2016). There are no universally accepted boundaries for this desert but it is expected to depend on stellar mass (McDonald et al., 2019). Adopting the conservative criterion of a period of 2 days (corresponding to irradiances 200), the “desert" has a single unambiguous denizen in our sample: KOI–952.05 (Kepler–32f), a member of a multi-transiting planet system (Swift et al., 2013), has an orbital period of 0.74 days and radius of , consistent with but more precise than that determined by Morton et al. (2016) and Berger et al. (2018) but significantly larger than previous estimates (e.g., Swift et al., 2013; Mann et al., 2013). There are five other Earth- and super-Earth-size Ultra Short Period (USP, d) planets in the sample. The smallest object in the sample is the candidate KOI–7793.01.
Cloutier & Menou (2020) found that the planet radius distribution hosted by stars with masses 0.08–0.42M⊙ (corresponding to spectral types M3-M8) differs from that of more massive stars in lacking a significant population of (sub)Neptunes. Ment & Charbonneau (2023) also found a very pronounced dearth of sub-Neptunes around mid-to-late type M dwarfs. Our sample does not substantially overlap in stellar mass with these analyses and falls off below 0.5, although there is at least one KOI hosted by a mid-M dwarf just above the radius valley (Fig. 3).
The location of the radius valley and its dependence on orbital period (a proxy for irradiance) is used to test models of escape of planetary atmospheres/envelopes (see Sec. 4). Following past work (e.g., Ho & Van Eylen, 2023, and references therein), we performed a regression on the full sample with support vector machines (SVM) as implemented in the Python sci-kit learn package, with a choice of regularization parameter of , following Van Eylen et al. (2019). Super-Earth and sub-Neptune classification was initially assigned initially based on radii smaller or larger than 1.68 (the value predicted for this sample from the relation in Ho et al. (2024)), then classification was iteratively redone (10 times) using the SVM. Confidence intervals were calculated with Monte Carlo simulations where planet radii were randomized according to their uncertainties. We plot the final assignments and converged regression () as the black line in Fig. 4, along with previously published regressions.
Our regression is only weakly period-independent and contrasts with the negative slopes for M dwarf host stars derived by Van Eylen et al. (2021, blue dashed line in Fig. 4), from combined solar- and late-type Kepler host stars derived by Ho et al. (2024) and evaluated at the 0.59M⊙ median sample mass of our sample (dashed magenta line), and for a sample of TESS-detected planets around M dwarfs with mass measurements (Bonfanti et al., 2024, dashed orange line). It is also inconsistent with the positive slope found by Cloutier & Menou (2020) for planets detected around late-type hosts by Kepler/K2 (dashed cyan line). However, the slope is consistent with that derived from a small sample of M dwarf planets with RV-determined masses by Luque & Pallé (2022). The cause of the disagreements between these estimates is not known, but could arise from sensitivity of the slope to objects with the shortest or longest orbital periods combined with refinement of radii by exclusion of binaries and improved stellar properties.
3.2 Corrected Radius Distribution
The detection efficiency-corrected radius distribution constructed for the entire small-planet sample is shown in the top panel of Fig. 5. The grey region spans the 68% confidence interval, and the individual planet radii are marked by the ticks at the bottom. The total occurrence is the integral over the distribution and is 1.84 planets per star. This presentation also shows the super-Earths and sub-Neptunes sa peaks separated by an intervening gap or “valley" around 1.6-1.7. We computed a valley depth, i.e. the ratio of the mean of the adjacent maximum to the minima, of 3.85. This is similar to that found for solar-type stars by Ho & Van Eylen (2023), and inconsistent with the trend of shallower valley with lower stellar mass found by Ho et al. (2024).
While Fig. 5 appears to resolve the super-Earth population into two overlapping peaks, this structure is not significant compared to the uncertainties in the reconstruction (represented by the shaded area). To verify this, for each of the 1000 bootstrap reconstructions we assessed whether the radius interval 1-1.7 contained one or two maxima and, in the latter case, calculated the ratio of the intervening minimum to the lowest of the two maxima. 75% of the reconstructions contain two maxima (at 1.2 and 1.4 ), but among these, the ratio of the minimum to lowest maximum is continuously distributed between about 0.6 to unity, with most values near unity. This indicates that the two observed peaks are a random occurrence from a continuous distribution of possible outcomes.
There are other, more ambiguous features in the radius distribution, including a “shoulder" at 3, representing four planets of similar size (Fig. 2). The radius distribution also hints at the existence of a population of Earth-size objects distinct from super-Earths (second peak from left in Fig. 5). This is not an artefact of detection bias, since the efficiency only decreases smoothly with radius in this regime (Fig. 2). The prominent peak at 0.64 in Fig. 5 is produced by KOI–314.03 (Kepler–138b), which has a low detection probability due to its comparatively wide (10.3 day) orbit. It is a member of a three-planet system and was discovered only after analysis with optimized, automated routines following the discovery of the other, larger members (Batalha et al., 2013). Its presence in the sample could be just a statistical fluke, or it could represent a large population of “sub-Earths" (Sinukoff et al., 2017).
3.3 Radius Evolution
We divided the planets with assigned ages into halves that are younger and older than the median gyrochronologic age of 3.86 Gyr (mean ages of 2.6 and 5.4 Gyr, respectively). These are plotted red and blue curves in the bottom panel of Fig. 5. The sub-Neptune peak is much less prominent in the older sample while the super-Earth is larger, with the ratio between the two (integrated over 1.68–4 and 0.7–1.68, respectively) declining from 2.38 to 0.58 from the younger to older moieties. This is consistent with the evolution of planets by the loss and/or contraction of low molecular-weight envelopes over Gyr. Indeed, the integrated probabilities over mini-Neptunes (1.7–4) fall from 0.59 to 0.31 and is roughly equivalent to the increase in the Earth- and super-Earth (0.7–1.7) from 0.25 to 0.53. There is no evidence for a decrease in the mean radius of the sub-Neptunes with age nor a deepening of the radius valley as has been seen in other analyses of solar-type stars (Petigura et al., 2022; Chen et al., 2022) but the valley is indistinct in the older sample due to the small number of sub-Neptunes. A two-sided Kolmogorov-Smirnov test provide weak statistical support for the young and old samples to be drawn from different populations ().
We investigated four systematic effects that could potentially mimic radius evolution: (1) an artefact due to a sensitivity of the radius distribution to the choice of divisor between younger and older sub-samples; (2) a variation in stellar properties with age that influences detection efficiency; (3) a difference in host star properties brought about by selecting for stars with detectable rotation period that in turn affects the detected planets; and (4) a variation in stellar properties (i.e. or mass) with age combined with a correlation between the planet population and stellar properties. We find no evidence that any of these is responsible for the observed difference in radius distribution between the younger and old sub-samples, and details are provided in Appendix A.
Figure 6 compares the period-radius distribution of the younger (red points) and older (blue points) sub-samples and hints at non-uniformity in this evolution. While there are both younger and older sub-Neptunes (, above the horizontal line) at periods days, nearly all those at longer periods (25 days) are younger. This is discussed further in Sec. 4.3.
4 Interpretation and Discussion
4.1 The Radius Valley
Models that explain the dichotomy between super-Earths and sub-Neptunes predict a dependence of the location of the radius valley on orbital period or stellar irradiation similar that can be compared to observations (e.g., see Table 6 in Ho & Van Eylen, 2023). Both photoevaporation and core-powered mass loss models predict negative slopes with period, in the former due to stellar X-ray and/or EUV radiation as the driver (e.g., Owen & Jackson, 2012), and in the latter a consequence of the role of stellar bolometric irradiation (Gupta & Schlichting, 2021). For example, Gupta & Schlichting (2019) derived an index of -0.11 for core-powered loss, independent of stellar mass, while Owen & Kollmeier (2017) derive -0.16 for photoevaporative loss. Such models are consistent with the slopes reported by Van Eylen et al. (2021), Ho et al. (2024), and Bonfanti et al. (2024) (Fig. 4). In contrast, a model of planet formation in a gas-depleted inner disk predicts a positive slope of +0.11 (Lopez & Rice, 2018), consistent with the findings of Cloutier & Menou (2020). None of these are consistent with the near-zero slopes found by Luque & Pallé (2022) and our analysis (; Fig. 4). Weak dependence on period as well as on stellar luminosity (Fig. 3), suggests a mechanism largely independent of the host star, e.g. differences in water abundance (see Burn et al., 2024, for a discussion of some possible scenarios).
The depth of the radius valley relative to the adjacent super-Earth and sub-Neptune peaks has also been compared to model predictions. In particular, Ho et al. (2024) found that the valley became shallower (filled in) at lower host star masses. A trend could appear due to error in the radius ratio from fitting the transit lightcurve, which, all else being equal, is expected to be higher for shorter duration transits around smaller M dwarf host stars (Petigura, 2020). Astrophysical interpretations of a shallower radius valley include: (a) scatter in the XUV-emission luminosity of young stars at a given mass and age; (b) heterogeneity in the composition of planetary cores; and (c) collisions (Ho et al., 2024). Our failure to reproduce the trend motivates a deeper dive into the systematics that could affect reconstructions of the planet radius distribution.
4.2 Evolution of Sub-Neptunes into Super-Earths
The decline of the sub-Neptune population and commensurate rise in super-Earths between our younger and older sub-samples (Fig. 5) indicates that some sub-Neptunes have evolved into super-Earths over a timescale of 3 Gyr. On orbits with day this radius evolution has presumably caused some planets to leave the sample due to the low detection efficiency of super-Earths on wider orbits (grey contours in Fig. 2). The observed change is analogous to that seen around solar-type Kepler host stars (Berger et al., 2020; Sandoval et al., 2021; David et al., 2021; Chen et al., 2022), which has been explained by a scenario where sub-Neptunes are super-Earths with thick hydrogen-dominated envelopes which can escape by XUV-driven photoevaporation or core-powered mass-loss from rocky/icy cores (Owen, 2019; Gupta & Schlichting, 2021; Owen & Schlichting, 2024).
High-energy emission from solar-type host stars is most intense during the first few hundred Myr (Johnstone et al., 2021), so a Gyr timescale for planet evolution has been previously interpreted as supporting core-powered (Berger et al., 2020). However, more quantitative estimates of photoevaporative timescales are longer (King & Wheatley, 2021) and both mechanisms may operate at different times in different regimes (Owen & Schlichting, 2024). Moreover, our sample cannot be used to investigate shorter timescales because it does not include systems with ages much younger than 1 Gyr. This is because low mass stars take longer to converge to an initial condition-independent rotation sequence, where gyrochronology can be applied.
In a scenario where sub-Neptunes are rocky cores with thick envelopes of high molecular weight volatiles, specifically H2O (Turbet et al., 2020), radius evolution could be at least partly explained by contraction due to cooling (Howe & Burrows, 2015; Vazan et al., 2018), and/or chemical reaction between the envelope and the rocky core (Kite et al., 2020; Kim et al., 2023). These processes have timescales different from that of atmospheric escape mechanisms.
4.3 The Cool Sub-Neptune Collapse
The near-complete absence of sub-Neptunes in the “desert" very close to the host stars is consistent with the canonical picture of envelope loss driven by stellar irradiation. However, the disappearance of sub-Neptunes elsewhere appears to be non-uniform (Fig. 6). There is a striking paucity of older sub-Neptunes on the widest orbits (20 days), where irradiation is lowest and escape should be slower than at intermediate orbits (2–20 days) where older sub-Neptunes are found (Fig. 6).
The top panel of Figure 7 shows planet radius versus stellar irradiance, comparing the younger and older sub-samples. Irradiance is based on bolometric luminosities calculated using interpolations of the corrections to -band luminosities tabulated in Pecaut & Mamajek (2013), and assuming near-circular orbits. Constructed irradiance distributions for younger and older sub-Neptunes (, dotted grey line) are shown in the bottom panel. Both Gaussian kernel density estimators (solid curves) and histograms with Poisson statistics errors (shaded bars) suggest a deficit of older sub-Neptunes at irradiances below several times the present terrestrial value (). The fraction of sub-Neptunes in the older-sub-sample falls from 53% at irradiances 5 to 17% at lower irradiance (see red and blue histograms). This cannot be due to a differences in detection efficiency since planets on wider orbits should be more readily detected around older, photometrically quieter stars. Nor is there evidence for a combination of correlations between age with host star mass (or luminosity), and stellar mass with planet period distribution (Figs. 14 and 15 in Appendix A) that could mimic this effect. There are no sub-Neptunes in the sample below an irradiance of 1, although this could be due in part to low detection efficiency at long orbital periods. The two planets that are marginally that are least irradiated are KOIs 812.03 (Kepler–235e) and 952.03 (Kepler–32d), with ages of and Gyr, respectively (Gaidos et al., 2023).
The persistence of “warm" sub-Neptunes and disappearance of cool sub-Neptunes at older ages could be due to an inflation mechanism that depends on stellar irradiation or planet equilibrium temperature and hence proximity to the host star. Turbet et al. (2020) showed that primarily rocky planets with water-dominated atmospheres could have sub-Neptune-like radii at irradiances exceeding a “runaway" threshold where H2O cannot condense into a liquid ocean.
Kopparapu et al. (2013) and Turbet et al. (2019) found the irradiance threshold for a runaway greenhouse on Earth-like planets to be 1.05. Around M dwarfs stars the threshold is slightly lower because of reduced Rayleigh scattering and increased absorption by atmospheric H2O at the longer wavelengths where cooler stars emit more radiation; the value is predicted to be 0.9 for =3900K or M0 spectral type (dashed vertical green line in Fig. 7. Three-dimensional general circulation models find a somewhat higher threshold that depends on planet rotation rate (Leconte et al., 2013; Kodama et al., 2018, ; vertical orange dashed line). The higher Bond albedo of a cloudy steam atmosphere (that of Venus is around 0.76 (Moroz et al., 1985) but measurements of the phase function suggest a value as high as 0.9 (Mallama et al., 2006)555The albedo depends would depend strongly on cloud droplet size and the high reflectivity of Venus is largely due to sulfuric acid aerosols and a similar scenario might not apply to steam atmospheres.), might move the runaway threshold to . This is the approximate boundary where sub-Neptunes become persistent (vertical dashed magenta line). Finally, atmospheres containing significant H/He will have a lower threshold (Innes et al., 2023).
While a decline in sub-Neptune occurrence across the water condensation boundary could, in principle, be explained by the collapse of steam atmospheres, it is unclear if H2O alone can explain the observed distributions with radius and age. In Fig. 7 we plot predictions of Turbet et al. (2020) for H2O mass fractions between 0.1 and 25% (cyan lines), assuming a 3 rocky core mass chosen to reproduce super-Earth radii. This comparison suggests that H2O mass fractions exceeding 25% are necessary to explain some sub-Neptunes. Models of H2O-rich planets produce radii up to 3 (Burn et al., 2024), but Vivien et al. (2022) found that pure water/steam atmospheres on sub-Earths become unstable if the atmospheric scale height exceeds 0.1 times the planet radius. Moreover, some internal governing mechanism (e.g., reaction with a magma ocean, Schaefer et al., 2016) is needed to explain the radius evolution over Gyr since the change in luminosity of these very low-mass host stars is small and the response of an atmosphere nearly instantaneous.
Alternatively, the inflated radii of sub-Neptunes and the differential evolution of cool vs. warm objects could be explained by atmospheres of mixed H/He and H2O. H/He lowers the mean molecular weight and decreases the irradiance threshold for a runaway water-vapor greenhouse (Innes et al., 2023). Slow escape of H/He would eventually collapse sub-Neptunes to smaller objects. At an elevated irradiances above the H2O condensation threshold, H2O could be persist in the upper atmosphere where its radiative cooling the lowers temperature and inhibits H/He escape (Yoshida et al., 2022). In a cooler sub-Neptune, water condenses into clouds deeper in the atmosphere and cannot play that role. The upper atmosphere is therefore hotter, H/He escape faster, and the sub-Neptunes eventually collapses to a rocky core with a geometrically thin atmosphere.
Finally, we searched for irradiance-dependent evolution of sub-Neptunes among Kepler transiting planets around solar-type host stars with isochrone ages determined by Berger et al. (2018), restricting the sample to those with ages more precise than 30%. Although the median age of the sample is 5 Gyr, we use the same division of 3.86 Gyr between young and old populations for comparison. No such effect is obvious in the irradiance distributions (Fig. 8). The ratio of young to old planets is statistically indistinguishable () for planets experiencing and . Three explanations for the difference between the two samples are (1) the late K/early M dwarf sample is affected by small-number statistics and is a statistical fluke; (2) the ages of the Berger et al. (2018) sub-sample with relatively cool planets are more uncertain because the host stars are likely to be late G and early K dwarfs which evolve very slowly in a color-magnitude diagram; and (3) there is a fundamental difference in the evolution of the two planet populations. At a given bolometric irradiance, planets around M dwarfs will experience far more X-ray and UV irradiation than their solar-type counterparts at (McDonald et al., 2019) and it is this radiation that drives photoevaporation of H/He atmospheres (Owen, 2019). Thus temperate sub-Neptunes around solar-type stars will have retained H/He atmospheres while their counterparts around M dwarfs have lost most or all their H/He and have radii controlled by the behavior of water-dominated atmospheres.
These different pathways are illustrated in Fig. 9, where the Kepler M dwarf sample spans Zones I, II, and III (but not Zone V due to small sample size), while the Berger et al. (2018) sample spans Zones I, II, and IV. In Zone I, rapid stellar-driven escape of H/He is not suppressed by atmospheric H2O and leads to planets with extended steam-rich atmospheres or rocky super-Earths (Fig. 9). In Zone II, water-rich (but not water-poor) sub-Neptunes retain thick H/He envelopes. In Zone III (only around very low-mass stars), H/He envelopes are lost, leaving “ocean worlds" and rocky super-Earths, and in contrast to Zones 4 and 5, where these envelopes are retained. Measurements of planet mass by RVs and observations of atmospheres by JWST for examples around much closer, brighter host stars could distinguish between these scenarios.
4.4 Caveats and Limitations of this Analysis
@@
Biases and selection effects that could possibly mimic the observed radius evolution are summarized in Sec. 3.3 and evaluated in Appendix A. There are other effects that could quantitatively impact our results. We restricted our sample to planet host stars with detected rotational signals in Kepler lightcurves and these are a biased subset of the overall population. Younger stars will be rapidly rotating, magnetically active, and highly spotted and their rotational periods can be readily measured. However, due that rotational variability as well as intense flaring can inhibit the detection of transiting planets (Gaidos et al., 2017). On the other hand, the oldest stars will be slowly rotating, less spotted, and exhibit less rotational variability, and the rotation periods can be similar to the Kepler quarterly roll ( days) and thus more difficult to retrieve (Santos et al., 2019). This is particularly true for the coolest stars in the sample (mid M-dwarfs) because these stars have the longest rotation periods for a given age. These selection effects can manifest themselves in the overall distribution of M dwarf planet host ages (e.g., Gaidos et al., 2023). However, we see no trend of planet properties (i.e., radius) with host star properties (i.e., luminosity, Fig. 3).
@@
Transit parameters and stellar parameters were not computed self-consistently: the former adopted DR25 values and the latter were based on revised luminosities and new empirical relations. Stellar density is covariant with transit impact parameter which, along with (which controls limb darkening coefficients) is covariant with the planet-to-star radius ratio . Lastly, our filtering of binaries from the Kepler late K- and early M-dwarf target sample was less thorough because AO imaging is not available for the overall Kepler late K/early M dwarf population as it is for the KOIs.
5 Summary and Future Directions
We refined the radii of 117 transiting planets detected by Kepler around 74 cool (3200–4200K) single dwarf stars and assigned ages to 113 based on their rotation and a revised gyrochronology. Based on these values, we conclude that: (1) analogous to the distribution found for solar-type host stars, the detection efficiency-corrected radius distribution of M dwarf planets contains two peaks representing super-Earths and sub-Neptunes, with the sub-Neptunes largely absent in a “desert" at periods days; (2) in contrast to some previous investigations of both solar-type and cooler dwarfs, the planet radius valley separating these two populations is only weakly dependent on orbital period (d/d); (3) there is a decline in sub-Neptunes and rise in super-Earths over the 3 Gyr separating the younger and older halves of the sample; and (4) the decline in sub-Neptunes is mostly at lower stellar irradiances where water can condense as clouds deep in the atmosphere. The last effect is not seen among Kepler planets around solar-type stars where the water condensation limit is much more distant. Taken together these suggest a governing role for intrinsic compositional differences in the radius dichotomy of small planets e.g., planets having water-rich and water-poor envelopes lying above and below the valley, respectively. The relative decline in cooler sub-Neptunes relative to their warmer counterparts with time could be due to cooling of the upper atmosphere by water and inhibition of escape of H/He.